a jet airplane is travelling at a speed of 500km/h.it ejects its products at the speed of 1500km/h relative to the jet plane.what is the speed of the latter with respect to an observer at the ground?

Grade:11

2 Answers

ali AskiitiansExpert-IIT
12 years ago

Dear Abhishek Choudhary,

      Velocity of jet w.r.t. ground Vjg= 500 km/h                                                                …...(Upward)

        Velocity of products relative to jet VPJ = 1500 km/h                                                …(downward)

Hence, velocity of  products relative to ground = -1500+500

                                                                                    =  -1000 km/h

Here – ve sing means downward .

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

Mr. Ali

IIT Delhi

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Velocity of jet w.r.t. ground Vjg= 500 km/h …...(Upward)
Velocity of products relative to jet VPJ = 1500 km/h ……(downward)
Hence, velocity of products relative to ground = -1500+500 = -1000 km/h
Here – ve sing means downward .

Thanks and Regards